### v5

# RPRO: Ranked Preference Reinforcement Optimization for Enhancing Medical QA and Diagnostic Reasoning

Abstract
Medical question answering requires advanced reasoning that integrates domain knowledge with logical inference. However, existing large language models (LLMs) often generate reasoning chains that lack factual accuracy and clinical reliability. We propose Ranked Preference Reinforcement Optimization (RPRO), a novel framework that combines reinforcement learning with preference-driven reasoning refinement to enhance clinical chain-of-thought (CoT) performance. RPRO distinguishes itself from prior approaches by employing task-adaptive reasoning templates and a probabilistic evaluation mechanism that aligns model outputs with established clinical workflows, while automatically identifying and correcting low-quality reasoning chains. Unlike traditional pairwise preference methods, RPRO introduces a groupwise ranking optimization based on the Bradley–Terry model and incorporates KL-divergence regularization for stable training. Experiments on PubMedQA, MedQA-USMLE, and a real-world clinical dataset from Far Eastern Memorial Hospital (FEMH) demonstrate consistent improvements over strong baselines. Remarkably, our 2B-parameter model outperforms much larger 7B–20B models, including medical-specialized variants. These findings demonstrate that combining preference optimization with quality-driven refinement provides a scalable and clinically grounded approach to building more reliable medical LLMs.
I Introduction
Recent advances in large language models (LLMs) signal significant promise for medical applications, especially in clinical reasoning and diagnostic support systems [singhal2023large, thirunavukarasu2023large, si2023evaluating, gilson2023does]. However, deploying AI in medicine requires exceptional reliability, making factual accuracy and thorough coverage of the data paramount. Medical reasoning introduces challenges distinct from those in general domains: clinical decisions require a comprehensive consideration of relevant factors, strict factual accuracy, and the avoidance of redundant information that could obscure critical insights.
Current approaches to medical reasoning primarily rely on chain-of-thought (CoT) prompting techniques [wei2022chain, kojima2022large, wang2023selfconsistency], which have shown effectiveness in improving reasoning capabilities across various domains. However, these methods typically generate single reasoning chains without systematic quality assessment or iterative refinement mechanisms. Moreover, existing evaluation frameworks often rely on overly simplistic binary metrics that fail to capture the nuanced quality required for medical reasoning or require extensive human annotation, limiting their suitability for automated quality control where domain expertise is essential.
Recent work in preference learning and reinforcement learning from human feedback (RLHF) [ouyang2022training, christiano2017deep, bai2022training] shows promise in aligning AI systems with human preferences. Direct Preference Optimization (DPO) [rafailov2023direct] has shown promising results by learning from preference data. However, these methods typically rely on pairwise comparisons, which fail to capture all the ranking information in human preference datasets. Many real-world scenarios involve humans ranking multiple candidates, not just providing binary choices. For example, when they evaluate reasoning chains, mathematical solutions, or creative content, evaluators assign rankings that reflect nuanced quality differences. Reducing this rich data to pairwise comparisons can lose valuable training signals and lead to suboptimal solutions.
Specifically, these approaches face significant challenges in medical domains. First, the requirement for extensive expert annotation is prohibitively expensive [talmor2020olmpics, joshi2023doctorsai]. Additionally, preference judgments often lack the granularity needed for medical reasoning assessment [lin2022human]. Furthermore, generic preference models fail to capture the domain-specific quality dimensions that are critical in healthcare applications [nori2023capabilities].
To address these limitations, we introduce a novel framework that enhances medical reasoning through probabilistic refinement for quality assessment and RPRO (Ranked Preference Reinforcement Optimization), which optimizes over ranked groups of candidates using the Bradley-Terry model. We combine domain-adaptive reasoning templates with automated quality assessment and targeted revision mechanisms. Unlike existing methods, which use single reasoning chains or expensive human evaluation, our framework automatically generates multiple reasoning candidates. It evaluates their quality using probabilistic metrics and performs targeted improvements when needed. Our approach incorporates several key innovations:
  • A probabilistic framework for automated medical reasoning quality assessment that captures multi-dimensional evaluation criteria through conditional probability modeling;
  • Domain-specific CoT enhancement templates that adapt reasoning structures to different medical contexts;
  • A novel linear reward learning framework that provides a stable policy gradient for ranked preference optimization;
  • A method for generating high-quality preference data suitable for Ranked Preference Reinforcement Optimization (RPRO) fine-tuning, enabling continual improvement of medical reasoning capabilities.
II Related Work
II-A Medical LLMs: QA, Diagnosis, and EHR
Large language models (LLMs) have been used for medical tasks like knowledge-based question answering, diagnostic reasoning on clinical vignettes, and patient-specific reasoning over electronic health records (EHR) [jin2019pubmedqa, jin2021medqa, shi2024ehragent, zakka2024almanac]. Domain-structured prompts and chain-of-thought (CoT) approaches consistently improve auditability and reduce unsupported claims by clearly separating decomposition and background from justification or diagnosis [wei2022cot, singhal2023medpalm, singhal2023medpalm2]. In patient-specific contexts, retrieval-augmented generation (RAG) and citation-style evidence tracing help ensure faithfulness to chart data and guidelines [lewis2020rag, shi2024ehragent, zakka2024almanac]. Recent work extends beyond answer accuracy, emphasizing multidimensional evaluation such as coverage of relevant findings and alignment with context, thereby better matching clinical expectations for completeness and correctness. In this work, we use domain-specific CoT templates (QA and diagnosis), automatically score rationales on coverage, factuality, and redundancy with an LLM judge, and make light score-guided revisions.
II-B Reinforcement-Learned Medical LLMs
Reinforcement learning from human feedback (RLHF) and related preference-based methods have become central to aligning LLM behavior beyond supervised fine-tuning [christiano2017deep, ziegler2019fine, ouyang2022training]. In reasoning-heavy domains, variants that regularize to a reference policy while optimizing rewards derived from preference ordering or task-specific outcomes have shown improved stability and faithfulness.
In the medical domain, reinforcement-learned LLMs have been applied to tasks such as clinical question answering and diagnostic reasoning [singhal2023towards, nori2023capabilities]. These methods typically employ reinforcement learning techniques such as policy optimization with preference feedback [christiano2017deep, ouyang2022training] to guide models toward multi-step clinical reasoning, integration of biomedical knowledge, and reliable elimination of alternative diagnoses. Nevertheless, prior studies also emphasize persistent challenges in maintaining stability and factual grounding in high-stakes clinical scenarios, indicating that preference-optimized training plays an important role in reducing hallucination risks [singhal2023towards].
Building on these advances, we introduce Ranked Preference Reinforcement Optimization (RPRO), a ranking-based extension of preference optimization. RPRO samples multiple responses per prompt and computes relative advantages, thereby reducing variance through group-level aggregation and providing more robust training signals for reasoning-intensive tasks like medical question answering.
II-C Ranking-based Optimization in Clinical Reasoning
Recent studies have expanded preference alignment for LLMs from pairwise to listwise settings. Most methods focus on pairwise comparisons. Yet, listwise approaches have shown effectiveness in other domains [liu2024lipo, pesaran2025lpoi, cai2025kpo]. For example, ALMupQA by [yang2024aligning] integrates multi-perspective ranking alignment for code QA. IRPO, introduced by [wu2025context], is a framework that incorporates graded relevance and positional importance. These works highlight the impact of ranking-based feedback on aligning LLM outputs with diverse user expectations. However, medical multiple-choice QA often needs more sophisticated explanatory discourse and clinical reasoning. In this work, we evaluate our framework on multiple medical reasoning benchmarks. We use a probabilistic refinement mechanism, generate preference data with RPRO, and combine linear rewards for ranked candidate optimization.
Figure 1: Overview of the proposed pipeline. The framework distinguishes between medical QA and diagnosis tasks, applies chain-of-thought (CoT) reasoning with multi-dimensional quality assessment and probabilistic refinement, and further improves through ranked preference reinforcement optimization (RPRO) with dataset construction.
III Methodology
III-A Overview
In this work, we provide an overview of a systematic framework for generating high-quality Chain-of-Thought (CoT) reasoning pairs optimized for Ranked Preference Reinforcement Optimization (RPRO) training in medical question answering. The framework follows a four-step pipeline (1) Based on medical task identification, we employ CoT reasoning for general medical Q&\&A and diagnostic tasks (2) For each chain instance, our process involves question decomposition, background knowledge integration, logical connection, and final justification for clinical assessment (3) We score each generated chain using probabilistic rules for assessment(4) After steps 1-3, we select high-quality reasoning chains and discard the lowest-quality onesThis results in a curated dataset for RPRO learning, enabling the model to learn from exemplary clinical reasoning patterns while avoiding reinforcement of suboptimal decision-making processes.
III-B Problem Formulation
We begin by formulating medical questions into two complementary task types to enable domain-appropriate reasoning. We define $\tau_{\text{QA}}$ as general medical question-answering, which targets questions requiring factual medical knowledge and logical reasoning about biomedical concepts, treatments, or underlying mechanisms. Conversely, $\tau_{\text{Diag}}$ represents medical diagnosis tasks involving clinical scenarios that necessitate systematic diagnostic reasoning through detailed patient presentation analysis and comprehensive differential diagnosis consideration. This dual-task framework establishes our task space $\mathcal{T}=\{\tau_{\text{QA}},\tau_{\text{Diag}}\}$. Let $\mathcal{X}$ and $\mathcal{Q}$ represent the spaces of structured medical records and natural-language prompts, respectively. Each task instance is an input pair $z=(x,q)\in\mathcal{X}\times\mathcal{Q}$, where $x$ is a medical record and $q$ is the corresponding prompt. A task classifier determines the appropriate task $\tau\in\mathcal{T}$ for each instance.
III-C Task-Adaptive Chain-of-Thought Generation
Given $z$ and $\tau$, a conditional policy $\pi_{\theta}(\cdot\mid z,\tau)$ generates $K=5$ candidate CoTs $C=\{c_{k}\}_{k=1}^{5}$ and $c_{k}\sim\pi_{\theta}(\cdot\mid z,\tau)$ for each medical question. Then, we present each candidate $c_{k}$ four reasoning chain steps as:
$c_{k}=\big(u^{(1)}_{k},\,u^{(2)}_{k},\,u^{(3)}_{k},\,u^{(4)}_{k}\big)$, (1)
where the $j$-th component $u^{(j)}_{k}$ depends on the task category $\tau$ and is defined as
$u^{(j)}_{k}=\begin{cases}s^{(j)}_{k},&\text{if }\tau=\tau_{QA},\\[6.0pt] t^{(j)}_{k},&\text{if }\tau=\tau_{Diag},\end{cases}\qquad j=1,2,3,4.$ (2)
For $\tau=\tau_{QA}$, the components include four clinical factors:
$\begin{array}[]{ll}s^{(1)}:&\text{Question Decomposition},\\ s^{(2)}:&\text{Background Knowledge},\\ s^{(3)}:&\text{Logical Connection},\\ s^{(4)}:&\text{Final Justification.}\end{array}$
to determine generated sequence medical QA chain is correct. Similarly, for $\tau=\tau_{Diag}$, which focuses on the diagnosis task, the components are defined as:
$\begin{array}[]{ll}t^{(1)}:&\text{Case Summary},\\ t^{(2)}:&\text{Clinical Significance},\\ t^{(3)}:&\text{Differential Diagnosis},\\ t^{(4)}:&\text{Most Likely Diagnosis.}\end{array}$
. Those criteria can better align the reasoning process with standard diagnostic procedures, thereby tailoring the CoT representation to the requirements of clinical decision-making. Thus, the unified formulation $c_{k}=(u^{(1)}_{k},u^{(2)}_{k},u^{(3)}_{k},u^{(4)}_{k})$ provides a task-agnostic representation of CoT candidates, while retaining the ability to specialize the semantic meaning of each reasoning step according to the task label $\tau$. After generation, the set of five candidates is reduced to $M=4$ by retaining the top four according to subsequent evaluation and discarding the lowest-ranked one.
III-D Medical Prompt Instruction and Reasoning
Previous research shows that effective prompt template instructions can improve model performance on CoT reasoning [zhu2024llms, wei2022chain, kojima2022large]. We use structured reasoning templates based on established clinical and biomedical reasoning patterns. Table V presents a general medical QA template that applies a four-step reasoning process, using binary classification at each step according to the validation criteria from Section III-C. Specifically, background knowledge retrieval gathers relevant biomedical facts and principles from medical literature. Logical connection links the retrieved knowledge to the question context. The justification step synthesizes the reasoning to support the conclusion.
In parallel, our diagnostic reasoning template mirrors the clinical diagnostic workflow used by medical professionals. The process begins with a case summary, which extracts and organizes key clinical findings from patient presentations. Clinical significance analysis then examines symptom patterns and diagnostic indicators to understand their medical relevance. Differential diagnosis systematically considers alternative diagnoses and establishes criteria for elimination. Finally, the most likely diagnosis step concludes with an evidence-supported diagnostic decision based on the available clinical information.
III-E Reasoning Refinement Method
III-E1 Multi-Dimensional Quality Assessment Evaluation
To evaluate reasoning quality across three key dimensions, we propose a joint-probability-based evaluation framework for assessing multiple candidate reasoning paths and selecting the optimal subset from $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$. For each candidate $c\in\mathcal{C}$, we define three scoring functions $S_{i}:\mathcal{C}\to\mathcal{S}$ where $i\in\{\text{cov},\text{fact},\text{red}\}$, corresponding to coverage $S_{\text{cov}}(c)$, factual accuracy $S_{\text{fact}}(c)$, and redundancy $S_{\text{red}}(c)$, each evaluated on a 5-point scale. These raw scores are mapped into probabilistic indicators $p_{i}(c)=\pi\big(S_{i}(c)\big)$ via a normalization mapping $\pi:\mathcal{S}\;\to\;[0,1]$.
III-E2 Probabilistic Reasoning Refinement
Unlike traditional additive scoring approaches, our objective is to simultaneously ensure adequate coverage of the source content and maintain factual accuracy, while also suppressing redundancy during generation. To achieve this, we formulate the decoding objective as the maximization of a multiplicative acceptance function that integrates these three criteria. Specifically, the each question of optimal candidate $c^{*}$ is defined as
$c^{*}=\arg\max_{c}\;\Big[\,p_{\text{cov}}(c)\cdot p_{\text{fact}}(c)\cdot\big(1-p_{\text{red}}(c)\big)\,\Big],$ (3)
where $p_{\text{cov}}(c)$ denotes the probability of coverage, $p_{\text{fact}}(c)$ denotes the probability of factual accuracy, and $p_{\text{red}}(c)$ measures the likelihood of redundancy in the generated content. Next, we introduce the selection operator with acceptance function $P_{\text{accept}}(c)$ as:
$\Phi_{k}(\mathcal{C})=\underset{c\in\mathcal{C}}{\operatorname{Top}\text{-}k}\;P_{\text{accept}}(c),$ (4)
where $\Phi_{k}(\mathcal{C})$ is extracts the top $k$ candidates with the highest acceptance probabilities to be used for subsequent training or decision-making. Intuitively, we can effectively unifies multi-dimensional quality assessment into a probabilistic formulation and provides a mechanism for ranking and optimal subset selection in reasoning tasks.
IV Training Framework
IV-A Ranking Preference Optimization Algorithm
In this section, we propose RPRO, a flexible optimization method that differs from traditional approaches such as DPO and PPO-RLHF, which rely solely on pairwise comparisons (i.e., $c_{1}\succ c_{2}$). Our proposed method leverages complete ranking information (i.e., $c_{1}\succ c_{2}\succ c_{3}\succ\ldots\succ c_{n}$) to optimize model performance. Our goal is to learn a policy $\pi_{\theta}(\cdot\mid z,\tau)$ that assigns higher likelihood to preferred responses. For each candidate CoT sequence $c_{j}$ given medical question $z$ and temperature $\tau$, we define the preference score as the conditional log-probability:
$s_{j}=\log\pi_{\theta}(c_{j}\mid z,\tau)=\frac{1}{|c_{j}|}\sum_{t=1}^{|c_{j}|}\log\pi_{\theta}(c_{j,t}\mid z,\tau,c_{j,<t})$ (5)
where $|c_{j}|$ denotes the sequence length of the CoT, and $c_{j,t}$ represents the token at position $t$ in the reasoning chain. Following the Bradley-Terry model [bradley1952rank], the probability that CoT sequence $c_{i}$ is preferred over sequence $c_{j}$ is:
$P(c_{i}\succ c_{j}\mid z,\tau)=\sigma\left(\frac{s_{i}-s_{j}}{\tau_{BT}}\right)=\frac{1}{1+\exp\left(-\frac{s_{i}-s_{j}}{\tau_{BT}}\right)}$ (6)
where $\tau_{BT}>0$ is a Bradley-Terry temperature parameter (distinct from the generation temperature $\tau$) controlling the sharpness of the preference distribution. Then, we can construct the ranking loss for the $K=5$ candidate CoT sequences extends the Bradley-Terry model to handle complete expert rankings. The $\mathcal{L}_{CoT-rank}$ can be rewritten as:
$\mathcal{L}_{CoT-rank}=\frac{1}{\binom{K}{2}}\sum_{i=1}^{K-1}\sum_{j=i+1}^{K}\log\left(1+\exp\left(-\frac{s_{i}-s_{j}}{\tau_{BT}}\right)\right).$ (7)
IV-B Linear Rewards From Rank
GRPO is one of the widely used algorithms that employs entropy regularization to stabilize policy gradient–based learning processes. However, it still exhibits sensitivity to parameter tuning, which can lead to unstable convergence performance or even divergence during training [leahy2022convergence, schulman2017proximal, mnih2016asynchronous]. To translate the complete ranking of candidate reasoning chains into effective training signals, we adopt a linear reward shaping scheme. We then define the advantage value $a_{j}$ of candidate $c_{j}$ as:
$a_{j}=(K-r_{j})-\frac{K-1}{2},\quad j=1,\dots,K\;\;\;\text{s.t.}\;\;\sum_{j=1}^{K}a_{j}=0.$ (8)
where each candidate is assigned a rank $r_{j}\in\{1,\dots,K\}$, with $r_{j}=1$ denoting the highest-quality candidate and $r_{j}=K$ the lowest. Intuitively, the linear reward can improve the model by incentivizing it to shift probability mass from poorly ranked reasoning paths toward higher-quality ones. Compared with pairwise-only preference optimization, RPRO efficiently incorporates full ranking information within each group of candidates, while retaining simplicity and stability during optimization. We then incorporate these linear rewards to increase the probability of higher-ranked candidates (positive $a_{j}$) and decrease the probability of lower-ranked ones (negative $a_{j}$) can be present as:
$\mathcal{L}_{\text{Linear}}=-\frac{1}{K}\sum_{j=1}^{K}a_{j}\,s_{j}.$ (9)
IV-C KL-Divergence Regularization for Medical Reasoning
To prevent the medical reasoning policy from deviating too far from a reference model $\pi_{ref}(\cdot\mid z,\tau)$, we incorporate KL-divergence regularization. The sequence-level KL divergence for CoT sequence $c_{j}$ is:
$\mathcal{D}_{KL}(c_{j}\mid z,\tau)=\frac{1}{|c_{j}|}\sum_{t=1}^{|c_{j}|}\mathcal{D}_{KL,t}$ (10)
The KL regularization term across all $K$ candidates is:
$\mathcal{L}_{KL}=\beta\cdot\frac{1}{K}\sum_{k=1}^{K}\mathcal{D}_{KL}(c_{k}\mid z,\tau)$ (11)
where $\beta\geq 0$ controls the regularization strength for medical reasoning stability. Here, the $\mathcal{D}_{KL,t}$ is token-level KL divergence at position $t$ in the CoT sequence be expressed as:
$\mathcal{D}_{KL,t}=\sum_{v\in\mathcal{V}}\pi_{\theta}(v\mid z,\tau,c_{<t})\log\frac{\pi_{\theta}(v\mid z,\tau,c_{<t})}{\pi_{ref}(v\mid z,\tau,c_{<t})}$ (12)
where $\mathcal{V}$ is the set of all possible tokens in language model. During training, we combine the $\mathcal{L}_{CoT-rank}$ and $\mathcal{L}_{KL}$ to form $\mathcal{L}_{RPRO-CoT}$ so that the final objective becomes:
$\mathcal{L}_{total}=\mathcal{L}_{RPRO-CoT}+\mathcal{L}_{\text{Linear}}$ (13)
V Experiments
V-A Experimental Setting
Datasets. We evaluate our approach on two benchmark medical QA datasets. PubMedQA [jin2019pubmedqa] contains biomedical research questions derived from PubMed abstracts, requiring yes/no/maybe answers based on scientific evidence. MedQA-USMLE [jin2021disease] consists of multiple-choice diagnostic questions from the United States Medical Licensing Examination, testing clinical reasoning and diagnostic capabilities. These datasets represent complementary aspects of medical reasoning: research-oriented knowledge synthesis and clinical diagnostic decision-making.
In addition, we further evaluate our model on the real-world Far Eastern Memorial Hospital (FEMH) clinical dataset, which consists of de-identified electronic medical records and physician diagnostic summaries. Unlike benchmark QA datasets, FEMH contains free-text clinical presentations and complex multi-symptom cases, providing a more realistic assessment of clinical reasoning performance.
Implementation Details. Our base model is Gemma 2B, chosen for computational efficiency while maintaining competitive performance. We generate five CoT reasoning variants per question and select the four highest-ranking candidates based on their acceptance probabilities for RPRO training. The acceptance threshold $\tau$ is set to 0.6 based on preliminary experiments. All models are trained with AdamW optimizer using a learning rate of 5e-5, batch size of 16, and 3 training epochs. Training is performed on NVIDIA A100 GPUs with mixed precision.
V-B Baselines
We compare against a diverse set of established language models spanning different parameter scales, training strategies, and levels of medical specialization. General-purpose LLMs include Gemma 7B [mesnard2024gemma], Mistral 7B [jiang2023mistral], Falcon 7B [penedo2023falcon], Qwen2.5 7B, and Vicuna 7B. These models represent strong modern baselines trained on broad web-scale corpora.
For medical-specialized models, we evaluate MedAlpaca 7B and 13B [han2023medalpaca], which are fine-tuned specifically for medical Q&A tasks, and BioMistral 7B [labrak2024biomistral], trained on biomedical literature and clinical text. We additionally include Medicine-LLM, a medical-domain model tailored for clinical reasoning tasks, and Meditron 7B [chen2023meditron], a strong medical adaptation of LLaMA trained on PubMed and clinical corpora. Finally, GPT-OSS 20B[agarwal2025gpt] is included as a large open-source frontier model with strong general and biomedical reasoning performance.
Together, these baselines cover a wide spectrum of model sizes 7B–20B), training regimes (general-purpose vs. biomedical-pretrained vs. medically fine-tuned), and architectural families, offering a comprehensive comparison against both domain-general and domain-specific state-of-the-art systems.
V-C Comparison with DPO and SFT
We compare our RPRO approach against two established training paradigms. Supervised Fine-Tuning (SFT) employs standard next-token prediction on the highest-quality reasoning chains. DPO utilizes pairwise preference learning between high- and low-quality reasoning pairs. RPRO implements group-wise ranking optimization across multiple quality levels, enabling more nuanced preference learning from ranked reasoning chains.
V-D Evaluation Metrics
We use two complementary metrics to assess model performance: accuracy, which measures the proportion of correctly answered questions for direct factual assessment, and macro F1 score, which computes the unweighted average of F1 scores across answer categories to ensure balanced evaluation and address class imbalance. This dual-metric approach captures both overall performance and category-specific reasoning quality.
TABLE I: Comparison of different models on PubMedQA and MedQA-USMLE under 0-shot, 1-shot, and 5-shot settings. Best results are in bold.
Models PubMedQA (Accuracy / Macro F1) MedQA-USMLE (Accuracy / Macro F1)
0-shot 1-shot 5-shot 0-shot 1-shot 5-shot
BioMedLM 35.10 / 20.45 42.23 / 25.75 48.65 / 30.30 22.68 / 13.73 26.07 / 15.92 31.18 / 20.36
BioMistral 7B 36.89 / 27.98 42.76 / 32.41 45.53 / 36.28 19.89 / 7.47 31.53 / 13.41 37.57 / 16.10
Falcon 7B 55.83 / 37.72 57.14 / 39.08 57.92 / 40.63 17.23 / 8.14 24.87 / 16.21 24.89 / 13.94
Gemma 7B 58.08 / 48.02 58.67 / 49.51 58.12 / 50.84 20.37 / 10.18 43.12 / 24.15 51.18 / 32.35
MedAlpaca 7B 44.83 / 42.41 46.72 / 43.94 48.51 / 45.34 24.01 / 16.37 41.27 / 31.46 47.82 / 36.38
Meditron 7B 55.32 / 44.82 58.65 / 43.50 60.44 / 46.27 19.23 / 14.77 27.88 / 19.94 34.01 / 24.33
Medicine-LLM 41.87 / 33.76 42.25 / 35.91 42.18 / 35.58 26.35 / 16.90 34.37 / 23.46 39.82 / 28.47
Mistral 7B 51.43 / 47.11 54.03 / 49.81 56.78 / 52.43 22.66 / 11.78 40.85 / 23.47 49.02 / 29.16
Qwen2.5 7B 51.34 / 48.40 54.43 / 46.82 55.38 / 43.44 35.06 / 23.84 44.23 / 31.02 50.87 / 37.68
Vicuna 7B 56.27 / 49.13 58.88 / 50.34 59.92 / 53.47 27.43 / 17.86 33.91 / 22.56 40.12 / 27.34
MedAlpaca 13B 52.61 / 46.16 52.47 / 47.93 54.38 / 49.29 23.15 / 22.42 39.36 / 25.27 45.31 / 26.54
GPT-OSS 20B 47.27 / 43.75 51.23 / 47.95 55.61 / 52.34 40.02 / 23.76 50.13 / 30.42 52.68 / 32.11
Our Method (Gemma 2B, RPRO) 60.23 / 51.78 61.76 / 50.95 62.02 / 53.12 42.38 / 30.51 46.27 / 32.68 51.67 / 35.24
TABLE II: Comparison of different models on the Far Eastern Memorial Hospital (FEMH) clinical dataset using BERTScore-F1 and Cosine Similarity under 0-shot, 1-shot, and 5-shot settings. Best results are in bold.
Models 0-shot (BERTScore-F1 / Cosine) 1-shot (BERTScore-F1 / Cosine) 5-shot (BERTScore-F1 / Cosine)
BioMedLM 0.798 / 0.342 0.812 / 0.351 0.825 / 0.353
BioMistral 7B 0.830 / 0.398 0.842 / 0.326 0.885 / 0.405
Falcon 7B 0.856 / 0.342 0.861 / 0.351 0.872 / 0.368
Gemma 7B 0.871 / 0.367 0.874 / 0.381 0.868 / 0.373
MedAlpaca 7B 0.848 / 0.334 0.852 / 0.343 0.861 / 0.357
Mistral 7B 0.859 / 0.345 0.865 / 0.359 0.871 / 0.366
Vicuna 7B 0.872 / 0.358 0.877 / 0.369 0.884 / 0.391
GPT-OSS 20B 0.851 / 0.338 0.864 / 0.354 0.873 / 0.368
Our Method (Gemma 2B, RPRO) 0.879 / 0.423 0.884 / 0.507 0.891 / 0.528
TABLE III: Ablation study comparing different training methods and refinement strategies on PubMedQA and MedQA-USMLE under zero-shot settings. Best results are in bold..
Method PubMedQA MedQA-USMLE
ACC (%) Macro F1 (%) ACC (%) Macro F1 (%)
SFT (with Refinement) 57.43 40.53 37.08 23.72
DPO (with Refinement) 59.90 42.72 38.26 24.35
RPRO (No Refinement) 59.29 37.41 37.53 23.47
Our Method: RPRO (with Refinement) 60.23 51.78 42.38 30.51
TABLE IV: Reasoning results from FEMH clinical dataset
[…]
Input (Symptoms/Objective):
S: 1101201 Enzyme decrease $\rightarrow$ possible improved myocarditis… 110/11/26 2nd Moderna $\rightarrow$ 11/27 fever $\rightarrow$ 11/29 mild dyspnea, chest tightness (worse when lying) … LMD CPK WNL, TnT 1559 → 210. EKG: mild pericarditis (r/o). CXR: normal heart size.
O: BP, HR stable; clear BS; soft abdomen; no edema.
Model Reasoning: Step 1. Identify the key symptoms: The patient presents with fever, dyspnea, chest tightness, and mild discomfort. Step 2. Consider the laboratory results: The patient has elevated troponin levels, but CK-MB is normal. This suggests a possible myocarditis. Step 3. Integrate the clinical findings: The patient’s symptoms and laboratory results suggest a possible myocarditis. The EKG is normal, and the CXR is normal. Step 4. Consider differential diagnoses: The differential diagnoses include myocarditis, pericarditis, and other causes of chest pain and dyspnea. Diagnosis: Myocarditis. Diagnosis: Myocarditis. Prediction & Answer:
Prediction: Myocarditis.
True Diagnosis: r/o mild myocarditis.
[…]
Input (Symptoms/Objective):
S: Stinging sensation (2021/11), high IOP (>>50), blurred vision OD. Lost f/u 2 years. O: IOP: 53→17 (treated). Cornea clear. PCIOL subluxation inferiorly. OS: NS 1+–2+. Fundus: C/D 0.5–0.6 (OD). OCT RNFL OD 32.
Model Reasoning: Step 1. Potential causes: corneal opacity, cataract, glaucoma, IOL subluxation. Step 2. History: high IOP, decreased vision, recent IOL subluxation. Step 3. Comprehensive exam to assess cornea/lens/retina. Step 4. Conclusion: likely IOL subluxation related to prior surgery/high IOP. Diagnosis: IOL subluxation. Diagnosis: IOL subluxation. Prediction & Answer:
Prediction: IOL subluxation.
True Diagnosis: IOL dislocation (OD).
[…]
Input (Symptoms/Objective):
S: Pain over lower right gingiva during upper RPD delivery.
O: Full-mouth chronic periodontitis, Upper/Lower RPD mostly acceptable except #46 supraocclusion, Ulceration on lower right edentulous ridge
Model Reasoning: Step 1. Assess the presence of an ulcer on the lower right edentulous ridge. Step 2. Determine if there is any retention or stability issue with the RPDs. Step 3. Evaluate the condition of the gingiva in relation to the RPDs. Step 4. Consider the history of chronic periodontal disease. Diagnosis: Chronic periodontal disease with ulceration. Prediction & Answer:
Prediction: Chronic periodontal disease with ulceration.
True Diagnosis: (1) Full-mouth chronic periodontitis. (2) Ulceration over lower right edentulous ridge mucosa.
Main Results. Table I presents the performance comparison across all evaluated models under 0-shot, 1-shot, and 5-shot settings on both datasets. Our method consistently achieves strong gains across all configurations. On PubMedQA, RPRO (Gemma 2B) attains 60.23%, 61.76%, and 62.02% accuracy, with corresponding Macro F1 scores of 51.78%, 50.95%, and 53.12% in the 0-shot, 1-shot, and 5-shot settings, respectively. Compared to the best baseline (Gemma 7B), this represents improvements of up to 3.9% in accuracy and 3.8% in Macro F1, demonstrating the robustness of our refinement mechanism across different few-shot conditions. On MedQA-USMLE, our model achieves 42.38%, 46.27%, and 51.67% accuracy, with Macro F1 scores of 30.51%, 32.68%, and 35.24%. While maintaining competitive accuracy, RPRO consistently delivers higher Macro F1 than the strongest baseline (GPT-OSS 20B), with gains of +6.75%, +2.26%, and +3.13% across the three settings, highlighting better reasoning completeness and answer quality. Overall, our 2B-parameter model matches or surpasses much larger 7B–20B models, demonstrating the effectiveness and efficiency of our reasoning refinement strategy.
Clinical Evaluation on FEMH Dataset. To further verify the generalization and clinical applicability of our approach, we conducted an evaluation on the real-world Far Eastern Memorial Hospital (FEMH) clinical dataset. Each case consists of de-identified electronic medical records containing a clinician-written subjective description (S) and objective examination findings (O), along with the ground-truth diagnostic assessment (A). The model is required to generate the diagnostic statement based solely on the S+O inputs, and the generated assessment is compared against the physician-annotated ground truth.
We evaluate the semantic similarity between the model-generated diagnosis and the true assessment using BERTScore-F1 and sentence-level Cosine Similarity under 0-shot, 1-shot, and 5-shot settings. As shown in Table II, our proposed RPRO (Gemma 2B) achieves the best performance across all configurations, obtaining BERTScore-F1 / Cosine scores of 0.879 / 0.423, 0.884 / 0.507, and 0.891 / 0.528 in the respective settings. Compared with larger medical-tuned models such as BioMistral-7B and MedAlpaca-7B, RPRO demonstrates higher semantic coherence and diagnostic consistency, highlighting strong few-shot generalization and reliable clinical reasoning despite its compact model size.
Representative reasoning examples generated by our method on the FEMH dataset are shown in Table IV. These cases illustrate how the model integrates subjective symptoms and objective findings to produce structured diagnostic reasoning that aligns closely with clinician-provided assessments.
Ablation Studies. We conduct a comprehensive set of ablation experiments to assess how different components and key hyperparameters influence the behavior of RPRO. In addition to examining the effects of various training objectives and refinement strategies, we focus on three critical hyperparameters: the optimization weight $\beta$, which controls the balance between supervised alignment and preference-driven updates; the rollout budget $K$, which determines the diversity and stability of sampled reasoning trajectories; and the acceptance threshold, which governs the selectivity of the refinement process. By systematically varying these factors across the PubMedQA, MedQA-USMLE, and FEMH datasets, we evaluate their impact on convergence behavior, reasoning quality, and overall task performance.
Ablation on Training Objectives and Refinement. We begin our analysis by examining how different training objectives and refinement strategies contribute to overall performance. As shown in Table III, standard supervised fine-tuning (SFT) with refinement provides a strong baseline, while DPO with refinement yields moderate improvements. RPRO without refinement already surpasses most baselines, demonstrating the advantage of groupwise ranking under the Bradley–Terry formulation. When refinement is added, RPRO achieves the best results across both PubMedQA and MedQA-USMLE, significantly improving both Accuracy and Macro F1. These results indicate that (1) the ranking-based preference optimization and (2) the reasoning refinement component are complementary and jointly responsible for the overall effectiveness of our method. This motivates the more fine-grained hyperparameter ablations presented in the following subsections.
Ablation on $\boldsymbol{\beta}$ Values. Fig. 2 presents the performance comparison under different $\beta$ values on the PubMedQA and MedQA-USMLE datasets. We vary $\beta\in\{0.0,0.05,0.1,0.2\}$ to control the balance between supervised alignment and preference-driven optimization. As shown in the figure, moderate $\beta$ values (around 0.1) yield the best results, achieving the highest Accuracy (60.23% on MedQA-USMLE, 42.38% on PubMedQA) and Macro F1 (55.83% and 30.51%, respectively). When $\beta$ is too small, the model underfits the preference signal; when $\beta$ is too large, over-regularization reduces reasoning diversity. This indicates that an appropriate trade-off between the base objective and preference regularization is crucial for stable optimization.
A similar trend is observed on the FEMH clinical dataset (right panel of Fig. 2), where $\beta=0.1$ achieves the best semantic alignment, with a BERTScore-F1 of 0.879 and a cosine similarity of 0.423. Both metrics degrade when $\beta$ deviates from this value, further confirming that moderate preference regularization leads to the most coherent and clinically meaningful diagnostic reasoning.
Figure 2: Performance comparison on PubMedQA, MedQA-USMLE, and FEMH under different $\beta$ values. PubMedQA/MedQA report Accuracy (solid blue) and Macro F1 (dashed orange), while FEMH reports BERTScore-F1 (solid blue) and Cosine Similarity (dashed orange).
Ablation on Rollout Budget $\boldsymbol{K}$. To further investigate the influence of rollout budget on reasoning quality, we vary the number of rollouts $K\in{2,4,5,8}$ during preference optimization. As illustrated in Fig. 3, increasing $K$ from 2 to 5 leads to steady improvements in both Accuracy and Macro F1 on PubMedQA and MedQA-USMLE. However, when $K$ exceeds 5, performance gains saturate and even slightly decline, indicating that excessive rollouts introduce redundant samples and reward noise, increasing computational cost without meaningful benefit. Empirically, K=5K{=}5 provides the best trade-off between performance and efficiency, consistent with prior observations in RLHF and DPO frameworks where higher rollout counts yield diminishing returns.
On the FEMH clinical dataset (right panel of Fig. 3), a similar trend is observed. Semantic similarity metrics including BERTScore-F1 and cosine similarity increase as $K$ grows to 4 or 5, reaching peak values of 0.879 and 0.423 (shown as 0.88 and 0.42 in the figure due to rounding). When $K$ is further increased, both metrics slightly degrade, suggesting that excessive rollouts introduce noisy or redundant chains of thought and weaken alignment. These results confirm that a moderate rollout budget yields the most coherent and clinically meaningful diagnostic reasoning.
Figure 3: Performance comparison on PubMedQA, MedQA-USMLE, and FEMH under different rollout numbers ($K$). PubMedQA/MedQA report Accuracy and Macro F1; FEMH reports BERTScore-F1 and Cosine Similarity. Each prompt generates $K$ candidate CoTs. For $K<4$, all candidates are used ($M=K$); for $K\geq 4$, the top $M=4$ candidates are selected for training.
Figure 4: Performance across different acceptance thresholds on PubMedQA, MedQA-USMLE, and FEMH. PubMedQA/MedQA report Accuracy (solid blue) and Macro F1 (dashed purple), while FEMH reports BERTScore-F1 and Cosine Similarity.
Threshold and Convergency Analysis. Fig. 4 examines the impact of the acceptance probability threshold $\tau$ on model performance. Lower thresholds (0.2, 0.4) result in excessive refinement, potentially degrading reasoning quality through over-correction. The optimal threshold of 0.6 balances the necessity for refinement with the preservation of reasoning. Higher thresholds (0.8) show marginal accuracy improvements but decreased Macro F1, suggesting that conservative refinement may miss opportunities for quality enhancement while maintaining correctness.
Figure 5: Training loss curves on MedQA-USMLE, PubMedQA, and FEMH. The plots show KL divergence, pairwise loss, ranking loss, and total loss across training steps for each dataset.
To evaluate the robustness of our training procedure, we further analyze the loss dynamics of RPRO as shown in Fig. 5. The results demonstrate that RPRO exhibits stable and convergent loss curves on both MedQA-USMLE and PubMedQA, confirming that the ranking-based preference optimization framework maintains reliable optimization behavior throughout training. The group-level ranking objective also effectively mitigates the high-variance fluctuations typically observed in pairwise preference learning.
On the FEMH clinical dataset (right panel of Fig. 5), we observe a similarly stable convergence pattern. Both the pairwise and ranking components decrease smoothly over training steps, and the total loss reaches a consistent downward trajectory without oscillation. This indicates that RPRO also achieves stable optimization when applied to real-world clinical narratives, further validating its robustness and applicability beyond benchmark QA datasets.
Reasoning Results from PubMedQA and MedQA-USMLE Tasks. The reasoning examples shown in Table V in Appendix A-B provide representative cases generated by our model using the proposed prompt templates. In PubMedQA, the model explicitly decomposes the question, incorporates relevant biomedical background knowledge, and connects evidence to the final answer, reflecting a literature-grounded reasoning pattern. In MedQA-USMLE, the reasoning traces follow a clinical workflow, including case summarization, identification of key findings, differential diagnosis, and treatment selection, demonstrating the model’s ability to mimic expert diagnostic procedures. Across both tasks, the generated reasoning showcases coherent multi-step inference, consistent structure, and appropriate use of domain knowledge. Together, these examples show that our approach yields structured and interpretable reasoning that is well aligned with expert biomedical and clinical decision-making, offering qualitative evidence of the model’s improved reasoning quality.
VI Conclusion and Future Work
In this paper, we introduced a novel framework for enhancing medical reasoning by integrating ranked preference reinforcement optimization with probabilistic reasoning refinement. Our method incorporates task-adaptive Chain-of-Thought generation, domain-specific reasoning templates, and a probabilistic quality assessment mechanism that enables targeted self-reflection and iterative improvement. Through evaluations on PubMedQA, MedQA-USMLE, and the FEMH clinical reasoning dataset, our 2B-parameter model demonstrates performance gains that surpass those of substantially larger baseline models. These findings highlight that quality-driven reasoning enhancement can be more impactful than parameter scaling alone. Furthermore, ablation studies validate the importance of reasoning refinement and ranked preference optimization in achieving consistent and reliable improvements across diverse medical tasks.
Looking ahead, we aim to extend our approach to a broader set of medical reasoning benchmarks and incorporate structured external medical knowledge to further improve factual grounding. Another promising direction is exploring optimal trade-offs between model scale, reasoning quality, and computational efficiency. Additionally, conducting human evaluation studies with clinical experts and real-world deployment assessments will be essential for understanding practical utility, reliability, and safety considerations. Ultimately, we envision this work contributing toward the development of trustworthy, high-quality AI systems for clinical decision support.
Acknowledgements
The authors would like to express their sincere gratitude to the FEMH AI Research Team for their invaluable support throughout the development of this work, including assistance with data preparation, system integration, and technical consultation. We are especially indebted to Dr. Fang-Ming Hung and Dr. Feng Liu for their continuous guidance, insightful feedback, and steadfast encouragement, which greatly enhanced the rigor and clinical relevance of this study. We also gratefully acknowledge the support from the Far Eastern Memorial Hospital Innovation Project, which provided essential resources and administrative coordination that enabled the successful execution of this research. This work was supported by the Far Eastern Memorial Hospital Innovation Research Project (Grant No. FEMH-2023-007).
VII Ethical approval
The study was approved by the Far Eastern Memorial Hospital (FEMH), Taiwan, Research Ethics Review Committee (FEMH IRB No.: 112086-F; https://www.femh-irb.org/ ), and all data were fully de-identified. All ethics review procedures and data collection were conducted in accordance with the committee’s standard guidelines and regulations (https://www.femh-irb.org/index.php/regulations ).
Conflict of interest
The authors declare that they have no competing interests or financial conflicts to disclose.
Appendix A Prompt Templates and Reasoning Outputs
A-A Prompt Templates
To improve the interpretability and reliability of model outputs, we introduce task-adaptive prompt templates that guide the generation of structured reasoning traces. Instead of relying on a one-size-fits-all format, each task type is paired with a reasoning scaffold that reflects its unique characteristics and objectives. This design choice serves two purposes: first, it enforces a step-by-step decomposition of the reasoning process, which makes the model’s decision pathway easier to trace and evaluate; second, it provides an inductive bias that steers the model towards domain-appropriate styles of justification.
The overarching principle behind these templates is to encourage multi-step reasoning that remains interpretable to human evaluators, while ensuring both factual accuracy and alignment with domain-specific norms. By providing explicit structures for reasoning, such as decomposition, background recall, logical connection, and justification, the templates reduce ambiguity in how the model organizes its answers. In turn, this structured approach facilitates both automated scoring and human assessment, making the reasoning process not only more transparent but also more consistent across different tasks.
General QA (PubMedQA-style)
This template is specifically tailored for biomedical question answering tasks such as PubMedQA, where the objective goes beyond producing a short answer. Instead, the emphasis is placed on generating an explicit reasoning trace that makes the answer both interpretable and verifiable. Biomedical questions often involve complex study designs, statistical results, or methodological considerations, and thus require reasoning that can bridge raw evidence with a clear conclusion.
By decomposing the reasoning into four stages: question decomposition, background knowledge, logical connection, and final justification, the prompt provides a structured scaffold that encourages the model to move progressively from understanding the question to constructing a justified answer. This explicit structure reduces the risk of shallow or incomplete responses and ensures that each step in the reasoning process contributes to the overall coherence of the final output.
Such a design is not only beneficial for improving factual accuracy, but also for enhancing transparency: human evaluators can more easily trace how the model arrived at its decision, identify possible reasoning gaps, and assess the quality of the justification. In this way, the prompt template serves as an inductive bias that steers the model toward producing consistent, domain-appropriate, and interpretable reasoning outputs.
Prompt Template (PubMedQA style) You are a medical reasoning expert. Based on the question, context, and final answer, produce a 4-step reasoning process: 1. Question Decomposition: Identify what the question is asking. 2. Background Knowledge: Recall relevant biomedical facts. 3. Logical Connection: Link knowledge to the question logically. 4. Final Justification: Conclude the answer with reasoning support. Ensure factual consistency, no hallucination.
Clinical QA (USMLE-style, as in MedQA)
For clinical question answering tasks that involve USMLE-style medical exam questions, such as those featured in the MedQA dataset, the prompt template is explicitly designed to emulate the diagnostic reasoning process of physicians. Unlike general QA, clinical cases demand not only factual recall but also structured interpretation of patient data, prioritization of differential diagnoses, and justification of the most plausible outcome. The reasoning scaffold is therefore organized into four stages: case summarization, where essential patient demographics and presenting symptoms are distilled; evaluation of clinical significance, which highlights the findings most relevant for decision-making; differential diagnosis, where possible explanations are systematically compared; and justification of the most likely diagnosis, which consolidates evidence in support of a single conclusion. By mirroring this stepwise diagnostic workflow, the template provides the model with an inductive bias towards producing reasoning that is aligned with clinical practice. This design facilitates both transparency and evaluability, enabling human experts to trace how the model arrives at its conclusions and to assess whether the reasoning follows medically sound logic.
Prompt Template (MedQA-USMLE style) You are a clinical reasoning expert. Given the patient case and final diagnosis, produce a 4-step diagnostic reasoning process: 1. Case Summary: Summarize the key patient information. 2. Clinical Significance: Explain the important findings. 3. Differential Diagnosis: Consider possible alternatives. 4. Most Likely Diagnosis: Justify the final diagnosis. Ensure medical accuracy, no hallucination.
A-B Model-Generated Reasoning Cases
The following reasoning cases, shown in Table V, present representative outputs generated by our model using the proposed prompt templates. These cases demonstrate how our approach enables structured, interpretable, and domain-specific reasoning across both biomedical and clinical question answering tasks.
Appendix B Scoring Examples
We illustrate how candidate CoTs are scored in our pipeline. Each candidate is systematically evaluated on three complementary criteria, using a 0–5 scale for each dimension: Coverage, which measures how comprehensively the reasoning process addresses the key aspects of the question; Factual Accuracy, which assesses whether the statements made are medically or scientifically correct; and Redundancy, which captures the extent of unnecessary repetition or irrelevant content that may obscure the clarity of the reasoning.
This tripartite scoring framework is designed to balance both the completeness and reliability of model-generated reasoning. By combining a positive dimension (coverage), a correctness dimension (accuracy), and a negative dimension (redundancy), we obtain a holistic view of quality that goes beyond whether the final answer is correct. In particular, the redundancy criterion ensures that verbose or circular explanations are not mistakenly rewarded, while the factual accuracy criterion penalizes hallucinations or incorrect claims even if the answer happens to be correct.
To make the evaluation process more concrete, we provide illustrative examples from two widely used benchmark datasets, PubMedQA and MedQA-USMLE. For each dataset, we present three representative CoTs: one high-quality, one medium-quality, and one low-quality. These examples are synthetic and included solely to demonstrate how our scoring framework operates, rather than being actual model outputs. Through these examples, it becomes clear that evaluating reasoning requires attention not only to the final answer but also to the intermediate steps that support it.
B-A PubMedQA
In PubMedQA, models are required to answer biomedical research questions and justify their predictions with evidence-based reasoning. To illustrate how our scoring framework operates, we present synthetic examples that highlight the distinction between comprehensive, factually grounded reasoning and superficial or vague chains of thought.
Example 1: High-quality CoT Question: Systematic use of patient-rated depression severity monitoring: is it helpful and feasible in clinical psychiatry? Answer: Yes Generated CoT: 1. Measurement-based care is key for achieving depression remission. 2. PHQ-9 was implemented in 19 psychiatric practices. 3. 93% of 6,096 patient contacts influenced by PHQ-9, 40% led to treatment changes. 4. Thus, PHQ-9 is both helpful and feasible in clinical psychiatry. Scores: Coverage = 4.7, Factual Accuracy = 4.9, Redundancy = 0.1
Example 2: Medium-quality CoT Question: Systematic use of patient-rated depression severity monitoring: is it helpful and feasible in clinical psychiatry? Answer: Yes Generated CoT: 1. PHQ-9 is sometimes used in psychiatry. 2. It measures depression symptoms. 3. Some patients changed treatment after using it. 4. Therefore, it is helpful. Scores: Coverage = 3.2, Factual Accuracy = 3.5, Redundancy = 0.6
Example 3: Low-quality CoT Question: Systematic use of patient-rated depression severity monitoring: is it helpful and feasible in clinical psychiatry? Answer: Yes Generated CoT: 1. Depression is bad. 2. Patients answer questions. 3. Doctors see results. 4. So it helps. Scores: Coverage = 1.2, Factual Accuracy = 1.5, Redundancy = 0.1
Figure 6: Examples of generated Chain-of-Thought (CoT) reasoning with varying quality levels on a PubMedQA sample. The high-quality CoT provides detailed, accurate, and non-redundant reasoning; the medium-quality example offers partial coverage; and the low-quality one demonstrates minimal reasoning coherence.
B-B MedQA-USMLE
For MedQA-USMLE, the task involves answering clinical questions in the USMLE style with precise diagnostic reasoning. To illustrate our scoring framework, we present examples showing that the same final answer can be supported by reasoning of very different quality. Our evaluation criteria capture whether the reasoning correctly contextualizes the case, enumerates differential diagnoses, and avoids vague or trivial statements.
Example 4: High-quality CoT Question: A 48-year-old man with chronic alcoholism presents with seizure, confusion, and severe hyponatremia. What is the most likely complication of urgent treatment? Answer: Osmotic myelinolysis Generated CoT: 1. Case Summary: 48M, seizure, confusion, history of alcoholism, hyponatremia. 2. Clinical Significance: Hyponatremia with seizure suggests electrolyte disturbance. 3. Differential Diagnosis: Possible complications include cerebral edema, arrhythmia, Wernicke encephalopathy. 4. Most Likely Diagnosis: Rapid correction of sodium can cause osmotic myelinolysis. Scores: Coverage = 4.8, Factual Accuracy = 5.0, Redundancy = 0.1
Example 5: Medium-quality CoT Question: A 48-year-old man with chronic alcoholism presents with seizure, confusion, and severe hyponatremia. What is the most likely complication of urgent treatment? Answer: Osmotic myelinolysis Generated CoT: 1. Patient had seizure and alcoholism. 2. Hyponatremia is dangerous. 3. Could be related to brain swelling or nerve problems. 4. So the answer is osmotic myelinolysis. Scores: Coverage = 3.0, Factual Accuracy = 3.2, Redundancy = 0.5
Example 6: Low-quality CoT Question: A 48-year-old man with chronic alcoholism presents with seizure, confusion, and severe hyponatremia. What is the most likely complication of urgent treatment? Answer: Osmotic myelinolysis Generated CoT: 1. Patient sick. 2. Sodium problem. 3. Bad outcome. 4. Answer: osmotic myelinolysis. Scores: Coverage = 1.0, Factual Accuracy = 1.5, Redundancy = 0.1
Figure 7: Examples of generated Chain-of-Thought (CoT) reasoning at different quality levels. Each example shows the same question and gold answer but varies in reasoning coverage, factual accuracy, and redundancy.
Appendix C Dataset Format (JSONL Example)
We adopt a JSONL format for our training data in the RPRO pipeline. Each entry corresponds to a single PubMedQA-style example, and contains the following fields:
  • id: unique identifier of the example.
  • question: the biomedical question to be answered.
  • context: supporting abstract or passage (from PubMedQA).
  • answer: the gold label.
  • ranked: a list of candidate reasoning chains ranked from better to worse.
  • scores: numerical quality scores corresponding to each reasoning chain.
  • breakdowns: fine-grained evaluation dimensions (coverage, factual accuracy, redundancy).
"id": "PMQA_001",
"question": "Can patient coaching reduce disparities?",
"context": "This study evaluated whether structured patient coaching improves adherence and reduces disparities ...",
"answer": "Yes",
"ranked": [
[…]
"scores": [
0.90,
0.65,
"breakdowns": [
{"coverage": 4.5, "factual_accuracy": 4.7, "redundancy": 0.2},
{"coverage": 3.5, "factual_accuracy": 3.2, "redundancy": 0.7},
Figure 8: An illustrative example from the PubMedQA dataset showing question, context, and ranked reasoning breakdowns with multi-dimensional quality scores.
TABLE V: Reasoning results for PubMedQA and MedQA-USMLE tasks.
Question Reasoning & Prediction
PubMedQA
Does body mass index (BMI) influence morbidity and long-term survival in gastric cancer patients after gastrectomy? Reasoning CoT:
1. Question Decomposition: Determine if BMI is a predictor of morbidity and long-term survival in gastric cancer patients.
2. Background Knowledge: BMI is a predictor of morbidity and long-term survival in gastric cancer patients.
3. Logical Connection: If BMI is significantly associated with these outcomes, it can serve as a predictive factor.
4. Final Justification: Based on the results of the study, BMI is a predictor of morbidity and long-term survival in gastric cancer patients.
Answer: yes
Final Prediction: yes
MedQA-USMLE
A 35-year-old man with itchy, watery eyes, recurrent in spring; conjunctival injection and watery discharge. Most appropriate treatment?
Options:
A. Erythromycin ointment
B. Ketotifen eye drops
C. Warm compresses
D. Fluorometholone eye drops
E. Latanoprost eye drops
Reasoning CoT:
1. Case Summary: 35M with itchy, watery eyes and sneezing, recurrent in spring, exam shows conjunctival injection with watery discharge.
2. Clinical Significance: These findings suggest allergic conjunctivitis rather than infection.
3. Differential Diagnosis: Bacterial conjunctivitis usually has purulent discharge, viral can be acute but less itchy, seasonal pattern fits allergy.
4. Most Likely Diagnosis: Allergic conjunctivitis, treatment is with antihistamine/mast cell stabilizer eye drops such as ketotifen.
Answer: Ketotifen eye drops (B)
Final Prediction = Ground Truth: B